1. Suppose the voltage rating of a batter is 180 V and the resistance are R1 R3-19 0 a What is the equivalent resistance? |2| = 3.20 , R2-780, b. What is the current through the circuit?


1
Expert's answer
2021-05-18T07:49:42-0400

1/Req= 1/(3.20)+ 1/(190)+1/(780)

=0.319043

Req=1/0.319043

=3.1"\Omega"

Current=V/R

=180÷3.1

=58.06"\Alpha"